OpenStudy (anonymous):

solve: 4096^(3x-4)=256

OpenStudy (anonymous):

u need help

OpenStudy (anonymous):

yea

Parth (parthkohli):

\[\left(2^{12}\right)^{\large 3x - 4} = 2^{8}\]

Parth (parthkohli):

If the bases are equal, then the powers will also be equal.

OpenStudy (anonymous):

x=14/9